OpenStudy (anonymous):

plzzz help find the angle between the vectors, approximate your answer to the nearest tenth: v=-5i-3j, w=2i-4j. A.85.6 degrees B.5.1 degrees C.147.5 degrees D.32.5 degrees

OpenStudy (misty1212):

hi take the dot product divide by the product of the absolute values take the inverse cosine you know how to get a dot product?

OpenStudy (misty1212):

it is very easy also very easy to cheat http://www.wolframalpha.com/input/?i=%28-5i-3j%29dot%282i-4j%29

OpenStudy (anonymous):

Can you take me through the steps plz

OpenStudy (misty1212):

but not really necessary \[-5\times 2+-3\times (-4)=-10+12=2\]

OpenStudy (misty1212):

you see how i got that ?

OpenStudy (anonymous):

yes

OpenStudy (misty1212):

ok now we need the absolute value of each one

OpenStudy (misty1212):

for the first one it is \[\sqrt{5^2+3^2}=\sqrt{25+9}=\sqrt{34}\]

OpenStudy (misty1212):

for the second one it is \[\sqrt{4^2+2^2}=\sqrt{16+4}=\sqrt{20}\]

OpenStudy (misty1212):

we can simplify those but no matter you are going to have to use a calculator anyways

OpenStudy (anonymous):

\[2\sqrt{5}\]

OpenStudy (misty1212):

looks like \(85.6^\circ\)

OpenStudy (anonymous):

i thought I was suppose to break down the 20

OpenStudy (misty1212):

like i said it doesn't really matter because you can’t do this without a calculator

OpenStudy (anonymous):

oooo ok thank you so much :)

OpenStudy (misty1212):

you are welcome, glad to help